Fermi National Laboratory
 
Operation Summary for November 30, 2011 - December 2, 2011
 

NuMI reported receiving 8.62E18 protons on target for the period from 11/28/11 to 12/5/11.

Friday December 2

The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with MiniBooNE off for air conditioning work at the MI-12 service building, and with the NuMI and Meson FTBFT-978 experiments taking beam.

During the day shift, Muon Rings personnel began a shift-long study. Booster personnel began a two-hour study at 8:54 AM. MI personnel began a MI-8 study at 10:41 AM. The MiniBooNE air conditioning work ended and MiniBooNE resumed taking beam at 3:32 PM. Operators reported short downtimes due to problems with three Booster RF stations (BRF17, BRF3, and BRF4).

During the evening shift, at 4:25 PM, a NuMI BPM (NBPM6N) stopped polling; operators rebooted the node. Operations disabled T-978 beam at 7:21 PM. At 9:23 PM, a Linac RF station (LRF5) tripped off; operators had to reset the station locally. At 10:05 PM, a M01 water skid
tripped off; operators called in a Water Group person. He reported that the problem was in the controller box. Operators called in an Instrumentation person.

Saturday December 3

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, the Instrumentation person reported that the Meson supply temperature was running a bit high; he raised the trip level at 12:38 AM. At 6:06 AM, the M01 water skid tripped off again; it reset. Operators adjusted the I- Source Arc power supply voltage at 7:14 AM. Operations established beam to T-978 at 7:48 AM, but this was only for a short time; problems with the M01 water skid held off beam while people investigated.

During the day shift, an Instrumentation person reported that he had raised the M01 temperature trip and made sure that the interlock would only trip on the supply temperature. Operations reestablished beam to T-978 at 9:51 AM.

During the evening shift, Operations disabled the T-978 beam at 7:04 PM.

Sunday December 4

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, Operations established beam to T-978 at 7:48 AM.

During the day shift, Operations loaded proton into the Recycler for studies at 8:14 AM. At 8:58 AM, operators reported that the I- Source output current was dropping and increased its output. Operators rebooted NBPM6N at 10:30 AM.

During the evening shift, at 4:26 PM, a Booster quadrupole magnet power supply tripped off on a ground fault indication; it reset. Operations disabled the T-978 beam at 7:01 PM.

Monday December 5

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, operators completed a 750 KeV line tune at 5:52 AM. Operators completed a Booster injection tune at 7 AM. Operations established beam to T-978 at 7:00 AM.

Operations Coordinator (OpsCo)

The OpsCo approved of the plans by Booster, Muon, and MI personnel to conduct studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/05/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory